# Kinetic Flat-Histogram Simulations of Non-Equilibrium Stochastic Processes with Continuous and Discontinuous Phase Transitions

### v1

@@ III Continuous Phase Transitions
We begin by applying the kinetic flat-histogram algorithm to several stochastic processes that exhibit continuous phase transitions. For all processes, we set the system size to $N=10^{3}$. The models considered are the Glauber model [43, 44, 45], the majority-vote (MV) model [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], the contact process (CP) [65, 66, 34, 67, 68, 69, 23, 70, 71, 24, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83], and the first Schlögl model [25, 26].
@@ III.1 Glauber Model
The Glauber model is a stochastic dynamics for the Ising model [43, 44, 45]. We note that the Glauber model obeys detailed balance, so we obtain the result for an equilibrium model in the asymptotic limit $t\to\infty$. The kinetic flat-histogram simulations also work for equilibrium systems. In the mean-field (one-site) limit, we can write the following reaction channels for the spin-flips:
@@
where $N=N_{\uparrow}+N_{\downarrow}$ is the system size. In Eq. (17), $k_{1}$ and $k_{2}$ are the spin-flip propensities:
@@
The magnetization $M$ obeys the following time-evolution equation:
@@
which predicts a continuous phase transition from a paramagnetic phase to a ferromagnetic one with an exact threshold $\lambda_{c}=1$. One can note that the stationary magnetization just obey the Bragg-Williams equation of state. The order parameter is the absolute value of the magnetization $m$, and the control parameter $\lambda$ acts as the inverse of the temperature.
Figure 1: (Color Online) Results for the Glauber model. In panel (a), we show the stationary distribution $P(M,\lambda)$ for a sequence of inverse temperatures $\lambda$. The $P(M,\lambda)$ has the expected behavior of a continuous phase transition, characterized by two symmetric maxima in the ferromagnetic phase ($\lambda>1$) and a single maximum in the paramagnetic phase ($\lambda<1$). In panel (b), we show a magnification of panel (a) where we emphasize the critical $P(M,\lambda_{c})$. The critical threshold $\lambda_{c}=1$ separates the two phases. Panel (c) shows the absolute value of the magnetization $m$. We see that the expected value from $P(M,\lambda)$ and the time-series average from SSA simulations are in good agreement. Finally, in panel (d) we show the Shannon entropy $S(\lambda)$, which is continuous. The line in panel (d) is only a guide to the eyes.
We show results for the kinetic flat-histogram algorithm in Fig. (1). The kinetic flat-histogram algorithm determines the stationary distribution $P(M,\lambda)$. $P(M,\lambda)$ exhibits the expected behavior for an inversion-symmetry breaking transition. Specifically, the critical $P(M,\lambda_{c})$ separates the paramagnetic phase, where the stationary distribution has a single maximum at $M=0$, from the ferromagnetic phase, where the stationary distribution has two symmetric maxima. There is good agreement between the results for the order parameter obtained using the kinetic flat-histogram simulations and the SSA algorithm. In addition, the continuity of the Shannon entropy is consistent with a continuous phase transition.
@@ III.2 Majority-Vote Model
Now, we consider the MV model [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64], which is a well-studied model of consensus formation. The MV model also describes a ferromagnetic material in contact with two heat reservoirs, one at $T=0$ and the other at $T\to\infty$. In the mean-field (one-site) regime, the model is equivalent to the Ehrenfest urn model.
@@
where the magnetization is still given by $M=(N_{\uparrow}-N_{\downarrow})/N$. The propensities mean that the spin opposes the majority direction with probability $\lambda$, follows the majority with probability $1-\lambda$, and has a random orientation if there is no majority.
The MV model magnetization obeys the following time-evolution equation:
@@
$\text{Sign}(x)=\begin{cases}-1&\text{if }x<0,\\ 0&\text{if }x=0,\\ 1&\text{if }x>0,\\ \end{cases}$ (22)
which predicts a continuous phase transition from a ferromagnetic phase to a paramagnetic one with an exact threshold $\lambda_{c}=0.5$. The order parameter is the magnetization $M$, and the control parameter acts as noise, defined as the probability that an individual opposes the opinion state of the majority.
Figure 2: (Color Online) Results for the MV model. The same as in Fig. (1), except that the system goes from the ferromagnetic phase to the paramagnetic phase by increasing $\lambda$, with a threshold $\lambda_{c}=0.5$.
We show results for the kinetic flat-histogram algorithm in Fig. (2). We again observe the expected behavior for an inversion-symmetry breaking transition, where the critical $P(M,\lambda_{c})$ separates the paramagnetic phase (with a single maximum at $M=0$) from the ferromagnetic phase (with two symmetric maxima). Similarly, the SSA and flat-histogram simulations are in good agreement. Furthermore, the continuous Shannon entropy is consistent with a continuous phase transition.
@@ III.3 Contact Process
The CP [65, 66, 34, 67, 68, 69, 23, 70, 71, 24, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] is a well-studied stochastic process that falls into the directed percolation universality class. It is the simplest stochastic process that presents an absorbing state. The reaction channels for the CP in the mean-field (one-site) regime are
@@
where $S$ denotes a susceptible individual and $I$ denotes an infected one. The first channel in Eq. (23) is an autocatalytic reaction representing contamination, with propensity $k_{1}=\lambda$. The second channel is spontaneous recovery, with propensity $k_{2}=1$. The dynamic behavior of the model is described by the infected concentration $\rho=N_{I}/N$, where the system size is $N=N_{S}+N_{I}$. Note that $\rho=0$ is an absorbing state.
The infected concentration obeys the following time-evolution equation:
@@
which predicts an active-absorbing phase transition with a threshold $\lambda_{c}=1$, and $\rho$ is the order parameter. The CP can describe epidemic spreading without permanent immunity, where $\lambda$ plays the role of the contamination probability and the recovery probability is unity. For $\lambda<1$, we have the absorbing phase where $\rho=0$, and for $\lambda>1$, we have the active phase where $\rho\neq 0$.
Figure 3: (Color Online) Results for the CP model. Panel (a) shows the stationary distribution $P(\rho,\lambda)$ for a sequence of contamination probabilities $\lambda$. For simplicity, we do not show $P(1,\lambda)$ and $P(2,\lambda)$ because these points are perturbed by the reactivation method [39]. The reactivation method is needed to make the CP dynamics ergodic, as the kinetic flat-histogram simulations require ergodicity. In the absorbing phase, $P(\rho,\lambda)$ presents a single maximum at the absorbing state $\rho=0$. In panel (b), are shown a magnification where we emphasize the critical $P(\rho,\lambda_{c})$ for the CP. The critical $P(\rho,\lambda_{c})$ approaches the absorbing state $\rho=0$ with a vanishing derivative. In panel (c), we can see that the expected value from Eq. (15) and the time-series average from the SSA algorithm are in good agreement. Finally, panel (d) shows the Shannon entropy $S(\lambda)$, which is continuous as expected for a continuous phase transition. The line in panel (d) is only a guide to the eyes.
We show results for the kinetic flat-histogram algorithm in Fig. (3). The absorbing-active phase transition exhibits a positive-definite order parameter. We note that the critical $P(\rho,\lambda_{c})$ separates the absorbing phase ($\lambda<1$) from the active phase ($\lambda>1$). In the absorbing phase, the stationary distribution $P(\rho,\lambda)$ shows a single maximum at the absorbing state $\rho=0$. In the active phase, $P(\rho,\lambda)$ shows a local minimum at the absorbing state. At the threshold, $P(\rho,\lambda)$ approaches the absorbing state with a vanishing derivative. Similarly, the continuous Shannon entropy is consistent with a continuous phase transition.
@@ III.4 First Schlögl Model
The first Schlögl model [25, 26] was introduced as a prototype of an open chemical reactive system exhibiting a continuous phase transition. It describes chemical reactions in rarefied gases with external particle sources, displaying a continuous non-equilibrium phase transition in the DP universality class. In the mean-field (one-site) regime, the reaction channels for the first Schlögl model are

### v2

@@ III Continuous Phase Transitions
We begin by applying the kinetic flat-histogram algorithm to several stochastic processes that exhibit continuous phase transitions. For all processes, we set the system size to $N=10^{3}$. The models considered are the Glauber model [20, 38, 56], the majority-vote (MV) model [16, 64, 85, 86, 13, 76, 55, 75, 28, 53, 54, 77, 44, 43, 5, 42, 78, 90, 2, 3], the contact process (CP) [33, 39, 35, 67, 59, 10, 36, 11, 17, 34, 8, 60, 25, 26, 62, 15, 24, 27, 4, 9, 21, 1], and the first Schlögl model [68, 31].
@@ III.1 Glauber Model
The Glauber model is a stochastic dynamics for the Ising model [20, 38, 56]. Notably, the Glauber model obeys detailed balance, so in the asymptotic limit $t\to\infty$. Indeed, the Glauber dynamics is a relaxation dynamics for the equilibrium Ising model. In the mean-field (one-site) limit, the spin-flip reaction channels can be written as
@@
where $N=N_{\uparrow}+N_{\downarrow}$ is the system size. In Eq. (17), $k_{1}$ and $k_{2}$ are the spin-flip propensities
@@
The magnetization $M$ obeys the following time-evolution equation
@@
which predicts a continuous phase transition from a paramagnetic phase to a ferromagnetic phase, with an exact threshold at $\lambda_{c}=1$. Note that the stationary magnetization satisfies the Bragg-Williams equation of state. The order parameter is the absolute value of the magnetization $m$, and the control parameter $\lambda$ acts as the inverse temperature.
Figure 1: (Color Online) Results for the Glauber model. Panel (a) shows the stationary distribution $P(M,\lambda)$ for a sequence of inverse temperatures $\lambda$. $P(M,\lambda)$ exhibits the expected behavior of a continuous phase transition, characterized by two symmetric maxima in the ferromagnetic phase ($\lambda>1$) and a single maximum in the paramagnetic phase ($\lambda<1$). Panel (b) presents a magnified view of panel (a), highlighting the critical $P(M,\lambda_{c})$. The critical threshold $\lambda_{c}=1$ separates the two phases. Panel (c) displays the absolute value of the magnetization $m$, demonstrating good agreement between the expected value from $P(M,\lambda)$ and the time-series average from SSA simulations. Finally, panel (d) shows the Shannon entropy $S(\lambda)$, which is continuous as expected for a continuous phase transition. The line in panel (d) is only a guide to the eye.
We present results for the kinetic flat-histogram algorithm in Fig. (1). The algorithm determines the stationary distribution $P(M,\lambda)$, which exhibits the expected behavior for an inversion-symmetry breaking transition. Specifically, the critical distribution $P(M,\lambda_{c})$ separates the paramagnetic phase, where the stationary distribution has a single maximum at $M=0$, from the ferromagnetic phase, where it has two symmetric maxima. There is good agreement between the order parameter values obtained using the kinetic flat-histogram simulations and those from the SSA algorithm. Additionally, the continuity of the Shannon entropy is consistent with a continuous phase transition.
Figure 2: (Color Online) Results for the Glauber model. Panel (a) shows the density of states $g(E)$ obtained using the original Wang-Landau algorithm. Panel (b) presents the stationary distribution $P(M,\lambda)$ calculated with both the kinetic flat-histogram algorithm and the original Wang-Landau algorithm, for 50 inverse temperatures $\lambda$ ranging from $0.8$ to $1.3$. The results are consistent, confirming the validity of the kinetic flat-histogram algorithm for equilibrium systems. However, the agreement deteriorates when going deep into the paramagnetic phase.
The Glauber model also allows us to test the kinetic flat-histogram algorithm for equilibrium systems, since the Glauber dynamics simulates the Ising model, whose Hamiltonian is given by
$\mathcal{H}=-\frac{J}{2N}\sum_{i,j}s_{i}s_{j},$ (20)
where $s_{i}=\pm 1$ are Ising spins and $J=1$ is the exchange coupling constant. From the density of states $g(E)$, one can calculate the stationary distribution $P(E,\lambda)$ in the canonical ensemble using
$P(E,\lambda)=\frac{g(E)e^{-\lambda E}}{Z(\lambda)},$ (21)
where
$E=-\frac{J}{2N}M^{2}$ (22)
is the mean-field energy and $Z(\lambda)$ is the partition function. We performed simulations using the original Wang-Landau algorithm to obtain $g(E)$ and calculated $P(M,\lambda)$ from it. The results agree with those obtained from the kinetic flat-histogram algorithm, confirming its validity for equilibrium systems, as shown in Fig. (2). The data from both simulations are consistent. However, the agreement deteriorates when going deep into the paramagnetic phase.
@@ III.2 Majority-Vote Model
We now consider the MV model [16, 64, 85, 86, 13, 76, 55, 75, 28, 53, 54, 77, 44, 43, 5, 42, 78, 90, 2, 3], a well-studied model of consensus formation. The MV model also describes a ferromagnetic material in contact with two heat reservoirs, one at $T=0$ and the other at $T\to\infty$. In the mean-field (one-site) regime, the model is equivalent to the Ehrenfest urn model.
@@
where the magnetization is still given by $M=(N_{\uparrow}-N_{\downarrow})/N$. The propensities indicate that the spin opposes the majority direction with probability $\lambda$, follows the majority with probability $1-\lambda$, and has a random orientation if there is no majority.
The magnetization $M$ in the MV model obeys the following time-evolution equation
@@
[…]
This equation predicts a continuous phase transition from a ferromagnetic phase to a paramagnetic one, with an exact threshold at $\lambda_{c}=0.5$. The order parameter is the magnetization $M$, and the control parameter $\lambda$ acts as noise, defined as the probability that an individual opposes the opinion state of the majority.
Figure 3: (Color Online) Results for the MV model. As in Fig. (1), the system transitions from the ferromagnetic phase to the paramagnetic phase as $\lambda$ increases, with a threshold at $\lambda_{c}=0.5$. For this case, we also present statistical error bars for the order parameter $m$ obtained from the kinetic flat-histogram simulations. The error bars, calculated from 10 independent runs, are smaller than the symbol size. The line in panel (d) is shown only as a guide to the eye.
We present results for the kinetic flat-histogram algorithm in Fig. (3). As expected for an inversion-symmetry breaking transition, the critical distribution $P(M,\lambda_{c})$ separates the paramagnetic phase (characterized by a single maximum at $M=0$) from the ferromagnetic phase (with two symmetric maxima). The order parameter values obtained from both SSA and flat-histogram simulations are in good agreement. Furthermore, the Shannon entropy remains continuous, consistent with a continuous phase transition. In Fig. (3), we also show statistical error bars for the order parameter $m$ obtained from 10 independent runs of the kinetic flat-histogram algorithm; these error bars are smaller than the symbol size.
To demonstrate the applicability of the method also to lattice systems, we present results for the MV model on square lattices in Fig. (4) and (5). New system states are generated using the kinetic Monte Carlo algorithm [45], where spin flips are attempted on randomly chosen nodes with rate
$w(\sigma_{i})=\frac{1}{2}\left[1-\left(1-2q\right)\sigma_{i}\text{Sign}\left(\sum_{\left<i,j\right>}\sigma_{j}\right)\right],$ (26)
with the summation performed over nearest neighbors and the sign function as previously defined. The local stochastic variable $\sigma_{i}=\pm 1$ represents the spin state of node $i$, and $q$ is the noise parameter. The kinetic Monte Carlo algorithm is equivalent to the SSA for lattice systems, and the time unit is defined as $N$ trials, where $N=L\times L$ is the system size.
Figure 4: (Color Online) Results for the MV model on a square lattice of size $N=50\times 50$. As in Fig. (1), the transition from the two-maxima regime to the single-maximum regime allows us to identify a pseudo-critical point $q^{\prime}=0.081$, which is close to the maximum of the Shannon entropy. The line in panel (d) is shown only as a guide to the eye.
Figure 5: (Color Online) Same as Fig. (4), except for $N=100\times 100$. We note that the pseudo-critical point $q^{\prime}=0.078$ approaches the critical threshold $q_{c}=0.075$ as the lattice size increases.
The spin-flip rate corresponds to the spin following the majority opinion in its neighborhood with probability 1q1-q, and the opposite with probability $q$, where $q$ acts as a social temperature by introducing local contrarians. After generating a trial move via kinetic Monte Carlo, the kinetic flat-histogram algorithm is applied by accepting trials with the probability given in Eq. (12). Alternatively, one may use the product of the spin-flip probability and the flat-histogram acceptance probability to generate system states. The results from kinetic Monte Carlo and flat-histogram simulations are in good agreement, as shown in Figs. (4) and (5).
The results for square lattices are qualitatively similar to those in the mean-field regime, except for finite-size effects. The transition from two maxima to one in the stationary distribution enables the detection of pseudo-critical points $q^{\prime}$ at noise values $q=0.081$ and $0.078$ for $L=50$ and $L=100$, respectively. The pseudo-critical point corresponds to the maximum of both the magnetic susceptibility and the Shannon entropy, and approaches the true critical point as the system size increases, following the scaling relation $q^{\prime}-q_{c}\sim L^{-1/\nu}$ [55, 78, 90]. Additionally, the continuous Shannon entropy is consistent with a continuous phase transition.
@@ III.3 Contact Process
The contact process (CP) [33, 39, 35, 67, 59, 10, 36, 11, 17, 34, 8, 60, 25, 26, 62, 15, 24, 27, 4, 9, 21, 1] is a well-studied stochastic process belonging to the directed percolation universality class. It is the simplest stochastic process that exhibits an absorbing state. In the mean-field (one-site) regime, the reaction channels for the CP are given by
@@
where $S$ denotes a susceptible individual and $I$ denotes an infected one. The first channel in Eq. (27) is an autocatalytic reaction representing contamination, with propensity $k_{1}=\lambda$. The second channel describes spontaneous recovery, with propensity $k_{2}=1$. The dynamic behavior of the model is characterized by the infected concentration $\rho=N_{I}/N$, where the system size is $N=N_{S}+N_{I}$. Note that $\rho=0$ is an absorbing state.
The infected concentration evolves according to the following time-evolution equation
@@
which predicts an active-absorbing phase transition with a threshold at $\lambda_{c}=1$, where $\rho$ serves as the order parameter. The CP can be used to describe epidemic spreading without permanent immunity, with $\lambda$ representing the contamination probability and the recovery probability set to unity. For $\lambda<1$, the system is in the absorbing phase with $\rho=0$, while for $\lambda>1$, it is in the active phase with $\rho\neq 0$.
Figure 6: (Color Online) Results for the CP model. Panel (a) shows the stationary distribution $P(\rho,\lambda)$ for a sequence of contamination probabilities $\lambda$. For simplicity, $P(1,\lambda)$ and $P(2,\lambda)$ are not shown, as these points are affected by the reactivation method [57]. The reactivation method is necessary to ensure ergodicity in the CP dynamics, which is required for kinetic flat-histogram simulations. In the absorbing phase, $P(\rho,\lambda)$ exhibits a single maximum at the absorbing state $\rho=0$. Panel (b) presents a magnified view, highlighting the critical $P(\rho,\lambda_{c})$ for the CP. The critical $P(\rho,\lambda_{c})$ approaches the absorbing state $\rho=0$ with a vanishing derivative. Panel (c) demonstrates good agreement between the expected value from Eq. (15) and the time-series average from the SSA algorithm. Finally, panel (d) shows the Shannon entropy $S(\lambda)$, which is continuous as expected for a continuous phase transition. The line in panel (d) is shown only as a guide to the eye.
We present results for the kinetic flat-histogram algorithm in Fig. (6). The absorbing-active phase transition is characterized by a positive-definite order parameter. The critical distribution $P(\rho,\lambda_{c})$ separates the absorbing phase ($\lambda<1$) from the active phase ($\lambda>1$). In the absorbing phase, the stationary distribution $P(\rho,\lambda)$ exhibits a single maximum at the absorbing state $\rho=0$. In the active phase, $P(\rho,\lambda)$ displays a local minimum at the absorbing state. At the threshold, $P(\rho,\lambda)$ approaches the absorbing state with a vanishing derivative. Additionally, the continuity of the Shannon entropy is consistent with a continuous phase transition.
@@ III.4 First Schlögl Model
The first Schlögl model [68, 31] was introduced as a prototype of an open chemical reactive system exhibiting a continuous phase transition. It describes chemical reactions in rarefied gases with external particle sources and displays a continuous non-equilibrium phase transition in the directed percolation (DP) universality class. In the mean-field (one-site) regime, the reaction channels for the first Schlögl model are written as
$A+B\xrightleftharpoons[k_{2}]{k_{1}}2A,\quad\text{and}\quad A\xrightleftharpoons[k_{4}]{k_{3}}C,$ (29)
which are combined with external sources and sinks for particles $B$ and $C$ that keep their concentrations constant in time, where $\rho=N_{A}/N$, $b=N_{B}/N$, $c=N_{C}/N$, and $N$ is the system size ($b$ and $c$ are constants). The dynamics are entirely described by the concentration $\rho$ of $A$ particles.
The concentration of $A$ particles obeys the following time-evolution equation
$\frac{d\rho}{dt}=-k_{2}\rho^{2}+(k_{1}b-k_{3})\rho+k_{4}c,$ (30)
where we set $k_{1}b=1$, $k_{2}=1$, and $k_{3}=k_{2}\lambda$. The system exhibits a continuous active-absorbing phase transition for $k_{4}=0$ as the control parameter $\lambda$ increases. Eq. (30), with these settings, predicts the threshold $\lambda_{c}=1$. For $k_{4}\neq 0$, there is no transition driven by the control parameter $\lambda$.
Figure 7: (Color Online) Results for the first Schlögl model. As in Fig. (6), the system is in the active phase for $\lambda<1$ and in the absorbing phase for $\lambda>1$.
Fig. 7 shows the stationary distribution $P(\rho,\lambda)$ obtained using the kinetic flat-histogram algorithm. The critical distribution $P(\rho,\lambda_{c})$ approaches the absorbing state with a vanishing derivative, similar to the behavior observed for the contact process. As in the active-absorbing phase transition of the contact process, the Shannon entropy remains continuous.
IV Discontinuous Phase Transitions
In this section, we present our findings for well-known bistable non-equilibrium stochastic systems in the mean-field (one-site) regime. Two of the most studied examples are the second Schlögl model [68, 31, 32, 82] and the Ziff-Gulari-Barshad (ZGB) model [89, 81, 18, 23, 35, 79]. These two stochastic processes provide a test to determine whether kinetic flat-histogram simulations can capture the physics of discontinuous phase transitions, especially weak ones. For all results, we set the system size to $N=10^{3}$.
IV.1 Second Schlögl Model
In parallel with the first Schlögl model, the second Schlögl model [68, 31, 32, 82] was introduced as a prototype of an open chemical reactive system exhibiting discontinuous phase transitions. It also describes chemical reactions in rarefied gases with external particle sources. In the mean-field (one-site) regime, the reaction channels for the second Schlögl model are given by
$2A+B\xrightleftharpoons[k_{2}]{k_{1}}3A,\quad\text{and}\quad A\xrightleftharpoons[k_{4}]{k_{3}}C.$ (31)
As in the first Schlögl model, particles $B$ and $C$ can be added or removed to keep their concentrations constant, where $\rho=N_{A}/N$, $b=N_{B}/N$, $c=N_{C}/N$, with $b$ and $c$ constants and $N$ the system size. Once again, the dynamics are entirely described by the concentration $\rho$ of $A$ particles.
The concentration of $A$ particles obeys the following time-evolution equation
$\frac{d\rho}{dt}=-k_{2}\rho^{3}+k_{1}b\rho^{2}-k_{3}\rho+k_{4}c,$ (32)
where we set $k_{1}b=1$, $k_{2}=1$, $k_{3}=k_{2}\lambda$, and $k_{4}c=\alpha$. Eq. (32), with these settings, predicts a line of discontinuous phase transitions with thresholds $\lambda_{\alpha}$. The critical line ends at a continuous phase transition with critical threshold $\lambda_{c}$ as $\alpha$ increases. When $\alpha=0$, we obtain the exact threshold of a discontinuous active-absorbing phase transition at $\lambda_{0}=1/4$, where the system is active for $\lambda<\lambda_{0}$ and absorbing otherwise. With our settings, the terminus of the critical line is the critical point at $\alpha_{c}=1/27$ and $\lambda_{c}=1/3$ [71].
Figure 8: (Color Online) Results for the second Schlögl model with $\alpha=0.005$. Panel (a) shows the stationary distribution $P(\rho,\lambda)$ for several values of the control parameter $\lambda$. We observe that $P(\rho,\lambda)$ exhibits two well-defined maxima as the threshold is approached. Panel (b) presents $P(\rho,\lambda)$ for two values of $\lambda$ close to the threshold $\lambda_{\alpha}$, one for each phase. Phase coexistence occurs when the two maxima of $P(\rho,\lambda)$ have equal weight. Panel (c) compares the expected value of the concentration $\rho$ with SSA simulations performed by increasing and decreasing $\lambda$. The SSA curves form a hysteresis cycle, which is a distinctive feature of discontinuous phase transitions. Finally, panel (d) shows the Shannon entropy $S(\lambda)$, which displays a clear discontinuity. The line in panel (d) is shown only as a guide to the eye.
We present results for the second Schlögl model in Figs. (8), (9), and (10), where $\alpha$ is increased as the terminus of the critical line is approached. For a strong discontinuous phase transition, as shown in Fig. (8), all the distinctive features of a discontinuous phase transition are evident. Even the SSA simulations reveal the hysteresis cycle. However, the hysteresis cycle is sensitive to the number of Monte Carlo steps (or the simulation time) discarded to allow the system to reach the stationary distribution. The stationary distribution displays two clear maxima, and the threshold can be estimated by identifying the $P(\rho,\lambda)$ corresponding to phase coexistence, when the two maxima have equal weight. Additionally, the Shannon entropy exhibits a clear jump near the threshold.
Figure 9: (Color Online) Results for the second Schlögl model with $\alpha=0.025$. As in Fig. (8), we are unable to observe two distinct maxima in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is clearly asymmetric, indicating a deviation from the symmetric parabolic shape. This provides evidence of a weak discontinuous phase transition, where the entropy jump is barely detectable. Additionally, the transition is too weak to exhibit a hysteresis cycle, and the Shannon entropy is practically continuous.
Fig. (9) presents results for a weak discontinuous phase transition. It becomes challenging to identify a discontinuous phase transition using standard methods. We discarded a sufficiently long initial simulation time to allow the system to reach the stationary distribution, so no hysteresis cycle is observed. Moreover, the jump in the Shannon entropy at the threshold begins to vanish. To observe clearer effects, larger system sizes should be used. Nevertheless, a distinctive feature remains: $\ln P(\rho,\lambda)$ is clearly asymmetric, allowing us to identify a possible weak discontinuous phase transition.
Figure 10: (Color Online) Results for the second Schlögl model with $\alpha=0.035$, close to the critical point at $\alpha_{c}=1/27\sim 0.037$ and $\lambda_{c}=1/3\sim 0.333$. As in Fig. (8), two distinct maxima are no longer visible in the stationary distribution $P(\rho,\lambda)$. However, $\ln P(\rho,\lambda)$ is nearly symmetric and close to a parabolic shape, indicating proximity to the critical point. Consequently, the transition is too weak to exhibit any hysteresis cycle, and the Shannon entropy is practically continuous.
Fig. 10 presents results near the critical point. It is difficult to observe any hysteresis cycle because importance sampling tunnels easily through phase space. Additionally, the Shannon entropy remains continuous. The only evidence of a discontinuous phase transition is the slight asymmetry of $\ln P(\rho,\lambda)$ near the maxima, indicating a minor deviation from the parabolic shape. If $\alpha$ is increased further, $\ln P(\rho,\lambda)$ near the maxima becomes parabolic and symmetric.
IV.2 Ziff-Gulari-Barshad Model
Another convenient test for the kinetic flat-histogram algorithm is the ZGB model [89, 81, 18, 23, 35, 79], which describes the catalytic oxidation of a metallic surface via the Langmuir-Hinshelwood process. The Langmuir-Hinshelwood process consists of three steps for a chemical reaction: (1) adsorption of the reactants onto a catalytic surface, (2) recombination of the reactants, and (3) desorption of the products.
The reaction channels of the ZGB model are given by
$C\xrightarrow{k_{1}}A,\quad 2C\xrightarrow{k_{2}}2B,$
$A+B\xrightarrow{k_{3}}2C,\quad\text{and}\quad A\xrightarrow{k_{4}}C,$ (33)
where $A$, $B$, and $C$ represent a carbon monoxide molecule, an oxygen atom, and an empty surface site, respectively. The channels from left to right in Eq. (33) are interpreted as follows:
  • First channel: adsorption of a carbon monoxide molecule, which occupies an empty site;
  • Second channel: adsorption of an oxygen molecule, where each of the two resulting atoms occupies an empty site;
  • Third channel: oxidation of carbon monoxide, resulting in a carbon dioxide molecule that is ejected from the surface, leaving two empty sites;
  • Last channel: possibility of a carbon monoxide molecule being ejected from the surface without recombination, leaving a single empty site.
This last channel allows for the existence of a critical line with a terminus at a critical point, as observed in the second Schlögl model.
Let $x_{1}$, $x_{2}$, and $x_{3}$ denote the concentrations of particles $A$, $B$, and $C$, respectively. The coupled time-evolution equations for the concentrations are given by